# Towards Wireless-Native Big AI Model: Insights into Its Ambitions, Peculiarities and Methodologies

### v2

@@ 5.5 Collaboration with AI-Agents
After constructing wBAIM, the next step is to consider the integration with AI-agent, an important booster for expanding the practical applications of current BAIM technology. Building upon wBAIM’s deep understanding of wireless regulations, adding modules such as planning, observation, memorization, and small model tool libraries will create wireless-native intelligent agent, significantly enhancing the usability and flexibility of wireless intelligence. For instance, a planning module can improve wBAIM’s logical reasoning capabilities and optimize its global performance; an observation module can enable autonomous acquisition of necessary state information, reducing signaling overhead; a memory module can help the agent to evolve continuously and reduce repetitive thinking; and small-tool libraries can lower the energy consumption required for intelligent computations and optimize the system configurations. An overall example is shown in Fig. 7. Furthermore, the technical crux of coordinating such multifunctional modules lies in controlling the overall computational latency and achieving effective fusion and alignment of multi-modal information.
@@ 6 Conclusion
Ubiquitous intelligence will become an indispensable and fundamental enabling technology as the next-generation wireless networks evolves towards more abundant functionality, more powerful performance and more flexible applications. To address the current challenges that wireless AI technologies face in terms of unit cost and reliability, BAIM, which has recently made significant breakthroughs in versatility and generalizability, have become a promising new approach. Therefore, how to build BAIM suitable for wireless networks has become a topic of great interest in wireless AI research.
In this study, we focus on the research roadmap of developing BAIM in a wireless-native way. We provide a detailed description of its scientific objectives and dissect the peculiarities and several informative methodologies in the evolution of wireless-native intelligence. We hope that this study will provide academic researchers and industry experts with some insightful ideas on the evolution toward wBAIM. Furthermore, we also look forward through further efforts in these research directions, wBAIM is anticipated to be built and bring a revolution in the intelligence of the whole wireless system.
Acknowledgement  This work was supported in part by National Natural Science Foundation of China under Grants 62394292, U20A20158 and 624B2129, Ministry of Industry and Information Technology under Grant TC220H07E, Zhejiang Provincial Key R&D Program under Grant 2023C01021, and the Fundamental Research Funds for the Central Universities No. 226-2024-00069.

### v3

@@ 5.5 Collaboration with AI-Agents
After constructing wBAIM, the next step is to consider the integration with AI-agent, an important booster for expanding the practical applications of current BAIM technology. Building upon wBAIM’s deep understanding of wireless regulations, adding modules such as planning, observation, memorization, and small model tool libraries will create wireless intelligent agent, significantly enhancing the usability and flexibility of wireless intelligence. For instance, a planning module can improve wBAIM’s logical reasoning capabilities and optimize its global performance; an observation module can enable autonomous acquisition of necessary state information, reducing signaling overhead; a memory module can help the agent to evolve continuously and reduce repetitive thinking; and small-tool libraries can lower the energy consumption required for intelligent computations and optimize the system configurations. An overall example is shown in Fig. 7. Furthermore, the technical crux of coordinating such multifunctional modules lies in controlling the overall computational latency and achieving effective fusion and alignment of multi-modal information.
@@ 6 Conclusion
Ubiquitous intelligence will become an indispensable, fundamental enabling technology as the next-generation wireless networks evolve and improve their functionality, performance, and flexibility. To address the current challenges that wireless AI technologies face in terms of unit cost and reliability, BAIM, which has recently made significant breakthroughs in versatility and generalizability, has become a promising new approach. Therefore, the question of building a BAIM suitable for wireless networks has become a topic of great interest in wireless AI research.
In this study, we focus on creating a research roadmap for developing native wBAIM. We provide a detailed description of its scientific objectives, dissect the peculiarities, and indicate instructive methodologies for the evolution of wireless intelligence. This study can provide academic researchers and industry experts with insightful ideas on the evolution of wBAIM. With further efforts in these research directions, wBAIM is anticipated to create a revolution in the intelligence of entire wireless systems.
Acknowledgement  This work was supported in part by National Natural Science Foundation of China under Grants 62394292 and 624B2129, Ministry of Industry and Information Technology under Grant TC220H07E, Zhejiang Provincial Key R&D Program under Grant 2023C01021, and the Fundamental Research Funds for the Central Universities No. 226-2024-00069.